A popular brand of cola contains 6.50g of carbon dioxide dissolved in 1 L of soft drink.  If the evaporating carbon dioxide is trapped in a cylinder at 1.00 atm and 20 degrees C, what volume does the gas occupy?
Answer #1
first,molar mass of CO2 is   12 + 16 + 16 = 44grams/mol
So calcthe number of moles:   6.50 grams * 1 mol / 44grams = 0.1477 moles
Now useideal gas law:
    p V = n R T    where T = 294K   and   p = 101,300 Pascals
Or:   V = nR T / p = 0.1477 * 8.31 * 294 / 101300 =0.00356286683m3 or 3.5628liters